### v1

# Optimal Control on Positive Cones

Abstract
An optimal control problem on finite-dimensional positive cones is stated. Under a critical assumption on the cone, the corresponding Bellman equation is satisfied by a linear function, which can be computed by convex optimization. A separate theorem relates the assumption on the cone to the existence of minimal elements in certain subsets of the dual cone. Three special cases are derived as examples. The first one, where the positive cone is the set of positive semi-definite matrices, reduces to standard linear quadratic control. The second one, where the positive cone is a polyhedron, reduces to a recent result on optimal control of positive systems. The third special case corresponds to linear quadratic control with additional structure, such as spatial invariance.
1 Introduction
The theory of optimal control has been a cornerstone of control theory ever since the early work by Bellman [1]. Under standard assumptions, a discrete time optimal control problem with non-negative cost
Minimize $\sum_{t=0}^{\infty}g(x(t),u(t))$
subject to $x(t+1)=f(x(t),u(t)),\quad x(0)=x_{0}$
$u(t)\in U(x(t))$
is solvable if and only if the Bellman equation
$J^{*}(x)$ $=\min_{u\in U(x)}\left[g(x,u)+J^{*}(f(x,u))\right]$
has a non-negative solution $J^{*}$ which is finite for all $x$. The minimizing $u$ gives an optimal control policy. See for example [2, section 3.1].
Of particular importance are the rare situations when the Bellman equation has an explicit solution. The most well known case is probably linear dynamics and quadratic cost, introduced in the pioneering work of [3]. Another case of fundamental importance is when both $x$ and $u$ are restricted to finite sets. This setting leads to shortest path problems on graphs, with an extensive literature dating back to the work of [4]. Recently, a third class of problems with explicit solutions to the Bellman equation was introduced in the context of linear systems on the positive orthant in $\mathbb{R}^{n}$, i.e. so called positive systems, see [5, 6, 7].
The purpose of this note is to show that the previous instances where the Bellman equation can be solved explicitly are in fact special cases of a more general class of optimal control problems stated in terms of positive cones.
2 Main Result
Theorem 1.
Let $\mathcal{P}$ be a positive cone in $\mathbb{R}^{n+m}$ with dual cone $\mathcal{P}^{*}$. Given $(s,r)\in\mathcal{P}^{*}$ and $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times m}$, suppose that for every $(x,u)\in\mathcal{P}$ there exists $v$ such that $(Ax+Bu,v)\in\mathcal{P}$. Moreover, assume the existence of $\phi:\mathbb{R}^{m}\to\mathbb{R}^{n}$ such that
$\phi(\mu)^{\top}x$ $=\min_{(x,u)\in\mathcal{P}}\mu^{\top}u$ (1)
for all $(x,v)\in\mathcal{P}$, $(\lambda,\mu)\in\mathcal{P}^{*}$. Then the following three statements are equivalent:
  • ($i$)
    The following optimal control problem has a finite value for every $(x_{0},u_{0})\in\mathcal{P}$:
    Minimize $\sum_{t=0}^{\infty}[s^{\top}x(t)+r^{\top}u(t)]\text{ over }\{u(t)\}_{t=0}^{\infty}$
    subject to $x(t+1)=Ax(t)+Bu(t)$
    $(x(t),u(t))\in\mathcal{P},\quad x(0)=x_{0}.$
  • ($ii$)
    There exists $(\lambda_{*},0)\in\mathcal{P}^{*}$ such that
    $\lambda_{*}=s+A^{\top}\lambda_{*}+\phi(r+B^{\top}\lambda_{*}).$ (2)
  • ($iii$)
    The convex program
    Maximize $\lambda^{\top}x_{0}\text{ over }(\lambda,0)\in\mathcal{P}^{*}$
    subject to $(s+A^{\top}\lambda-\lambda,r+B^{\top}\lambda)\in\mathcal{P}^{*}.$
    has a bounded value for every $(x_{0},u_{0})\in\mathcal{P}$.
If ($ii$) is true, then the minimal value of ($i$) and the maximal value of ($iii$) are both equal to $\lambda_{*}^{\top}x_{0}$. Moreover, the optimal control law $x\mapsto u$ is obtained by minimizing $(r+B^{\top}\lambda_{*})^{\top}u$ over $u$ subject to the constraint $(x,u)\in\mathcal{P}$.
Proof.
The minimization in ($i$) is a special case of the general formulation in the introduction, where $f(x,u)=Ax+Bu$ and $g(x,u)=s^{\top}x+r^{\top}u$. Define the sequence $\{\lambda_{k}\}_{k=0}^{\infty}$ recursively by
$\lambda_{k+1}$ $:=s+A^{\top}\lambda_{k}+\phi(r+B^{\top}\lambda_{k}),$ $\lambda_{0}$ $:=0.$
If $J_{k}(x)=\lambda_{k}^{\top}x$, then
$\min_{u\in U(x)}\left[g(x,u)+J_{k}(f(x,u))\right]$
$=\min_{(x,u)\in\mathcal{P}}\left[s^{\top}x+r^{\top}u+\lambda_{k}^{\top}(Ax+Bu)\right]$
$=s^{\top}x+\lambda_{k}^{\top}Ax+\min_{(x,u)\in\mathcal{P}}\left[(r+B^{\top}\lambda_{k})^{\top}u\right]$
$=\left[s+A^{\top}\lambda_{k}+\phi(r+B^{\top}\lambda_{k})\right]^{\top}x$
$=\lambda_{k+1}^{\top}x.$
Hence, value iteration
$J_{k+1}(x)$ $=\min_{u\in U(x)}\left[g(x,u)+J_{k}(f(x,u))\right],$ $J_{0}$ $=0$
gives a sequence $0=J_{0}\leq J_{1}\leq J_{2}\leq\ldots$ with $J_{k}(x)=\lambda_{k}^{\top}x$ for all $k$. In particular $(\lambda_{k},0)\in\mathcal{P}_{*}$ for all $k$. If ($i$) holds, then the sequence has an upper limit $J^{*}(x)=\lambda_{*}^{\top}x$ with $\lambda_{*}$ satisfying the condition in ($ii$). Conversely, if ($ii$) holds, then $(\lambda_{*}-\lambda_{k},0)\in\mathcal{P}_{*}$ for all $k$, so the sequence $J_{0}\leq J_{1}\leq J_{2}\leq\ldots$ has an upper limit and ($i$) holds too. This proves the equivalence of ($i$) and ($ii$), as well as the formula $\lambda_{*}^{\top}x_{0}$ for the minimal value in ($i$). The optimal controller is $u(t)=\lambda(x(t)$, where
$\lambda(x)$ $=\arg\min_{(x,u)\in\mathcal{P}}\left[s^{\top}x+r^{\top}u+\lambda_{*}^{\top}(Ax+Bu)\right]$
$=\arg\min_{(x,u)\in\mathcal{P}}\left[(r+B^{\top}\lambda_{*})^{\top}u\right].$
It remains to prove equivalence betwen ($ii$) and ($iii$).
First, by definition of $\phi$, the optimization constraint in ($iii$) implies that
$\lambda^{\top}x$ $\leq s^{\top}x+r^{\top}u+\lambda^{\top}(Ax+Bu)$ (3)
for all $(x,u)\in\mathcal{P}$. It is a well known fact that solutions to the Bellman inequality are lower bounds on the optimal cost. Hence, if ($ii$) holds, the maximum for the convex program must be attained at the solution of eq. 2, so also ($iii$) holds. Conversely, suppose that ($iii$) holds with the maximizing argument $\lambda_{k}$. Define
$\lambda_{k+1}$ $:=s+A^{\top}\lambda_{k}+\phi(r+B^{\top}\lambda_{k}).$
It follows from eq. 3 that
$\lambda_{k}^{\top}x$ $\leq s^{\top}x+r^{\top}u+\lambda_{k}^{\top}(Ax+Bu)$
$\lambda_{k}^{\top}x$ $\leq s^{\top}x+\lambda_{k}^{\top}Ax+\phi(r+B^{\top}\lambda_{k})^{\top}x$
for all $(x,u)\in\mathcal{P}$ and therefore
$(\lambda_{k+1}-\lambda_{k},0)$ $=(s+A^{\top}\lambda_{k}+\phi(r+B^{\top}\lambda_{k})-\lambda_{k},0)\in\mathcal{P}_{*}.$
It follows that
$\lambda_{k+1}^{\top}x$ $\leq s^{\top}x+r^{\top}u+\lambda_{k+1}^{\top}(Ax+Bu)$
Hence also $\lambda_{k+1}$ satisfies the constraints of the optimization in ($iii$), so $\lambda_{k+1}=\lambda_{k}$. Otherwise $\lambda_{k}$ would not be maximizing. This means that $\lambda_{k}$ solves eq. 2 and also ($ii$) holds. The proof is complete. ∎
1 supposes the existence of a function $\phi$ with the key property that it renders the optimization in eq. 1 linear in $x$. It is this feature that keeps the value iteration linear in $x$, which is instrumental in allowing the recasting of the optimal control problem as either a fixed point equation or convex program (conditions ($ii$) and ($iii$) in 1). It is natural then to ask when such a $\phi$ exists. The following result shows that when the cone $\mathcal{P}$ is proper, this can be established in a rather direct way through a property of the dual cone $\mathcal{P}^{*}$. We will use this result in the next section when we connect 1 to some well studied optimal control problems.
Theorem 2.
Let $\mathcal{P}$ be a proper cone in $\mathbb{R}^{n+m}$ with dual cone $\mathcal{P}^{*}$, $\lambda^{*}\in\mathbb{R}^{n}$, and $\mu\in\mathbb{R}^{m}$. Suppose that there exists a $\lambda\in\mathbb{R}^{n}$ such that
$\left(\lambda,\mu\right)\in\operatorname{rel}\operatorname{int}\mathcal{P}^{*}.$ (4)
Then the following statements are equivalent:
  1. ($i$)
    $-\lambda^{*\top{}}x=\inf_{\left(x,u\right)\in\mathcal{P}}\mu^{\top}u.$ (5)
  2. ($ii$)
    $\left(\lambda^{*},\mu\right)\in\mathcal{P}^{*},\;\text{and for all}\;\left(\lambda,\mu\right)\in\mathcal{P}^{*},\;\left(\lambda-\lambda^{*},0\right)\in\mathcal{P}^{*}.$
Proof.
($i$$\implies$($ii$): With a view to generating a contradiction, suppose that under the hypothesis of ($i$),
$\left(\lambda^{*},\mu\right)\notin\mathcal{P}^{*}.$
There therefore exists an
$\left(\bar{x},\bar{u}\right)\in\mathcal{P}\;\text{such that}\;\begin{bmatrix}\lambda^{*}\\ \mu\end{bmatrix}^{\top}\begin{bmatrix}\bar{x}\\ \bar{u}\end{bmatrix}<0,$
which implies that $-\lambda^{*\top{}}\bar{x}>\mu^{\top}\bar{u}$. However from eq. 5 we see that $-\lambda^{*\top{}}\bar{x}\leq{}\mu^{\top}\bar{u}$, which is a contradiction. This proves the first claim in ($ii$). Now, again with a view to generating a contradiction, suppose that
$\left(\lambda,\mu\right)\in\mathcal{P}^{*}\;\text{and}\;\left(\lambda-\lambda^{*},0\right)\notin\mathcal{P}^{*}.$
Therefore there exists an
$\left(\tilde{x},\tilde{u}\right)\in\mathcal{P}\;\text{such that}\;\begin{bmatrix}\lambda-\lambda^{*}\\ 0\end{bmatrix}^{\top}\begin{bmatrix}\tilde{x}\\ \tilde{u}\end{bmatrix}<0,$
which implies that $\lambda^{*\top{}}\tilde{x}>\lambda^{\top}\tilde{x}$. The dual of eq. 5 reads
$\sup_{\left(\lambda,\mu\right)\in\mathcal{P}^{*}}-\lambda^{\top}x.$ (6)
The solution to eq. 6 is lower bounded by $-\lambda^{\top}\tilde{x}$. However by weak duality, the solution to eq. 6 is upper bounded by $-\lambda^{*\top{}}\tilde{x}$. Therefore
$\left(\lambda,\mu\right)\in\mathcal{P}^{*}\implies\lambda^{*\top{}}\tilde{x}\leq{}\lambda^{\top}\tilde{x}.$
Again this a contradiction, proving the claim.
($ii$$\implies$($i$): Since $\mathcal{P}$ is a proper cone, $\mathcal{P}^{**}=\mathcal{P}$. Therefore for all $\left(x,u\right)\in\mathcal{P}$,
$\begin{bmatrix}\lambda-\lambda^{*}\\ 0\end{bmatrix}^{\top}\begin{bmatrix}x\\ u\end{bmatrix}\geq{}0,\implies-\left(\lambda-\lambda^{*}\right)^{\top}x\leq{}0.$
Therefore for any such $\lambda$,
$-\lambda^{\top}x=-\left(\lambda-\lambda^{*}\right)^{\top}x-\lambda^{*\top{}}x\leq{}-\lambda^{*\top{}}x.$
It then follows that the solution to eq. 6 equals $-\lambda^{*\top{}}x$. Condition eq. 4 states that eq. 6 is strictly feasible. This implies that strong duality holds, which proves eq. 5 as required. ∎
3 Examples
3.1 The Linear Quadratic Regulator
In this example we will connect the standard linear-quadratic-regulator (LQR) problem to 1. This will give an interpretation of our more abstract notation and results in a hopefully more familiar setting. In particular we will show that when the optimization in ($i$) corresponds to the LQR problem, conditions ($ii$) and ($iii$) reduce to the algebraic Riccati equations and semi-definite programs that are standard in the solution to this problem.
We will in fact study a slightly more general problem. The connections to the standard LQR will be established at the end of the section. The optimization problem that we consider first is:
Minimize $\sum_{t=0}^{\infty}\operatorname{tr}\left(\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\right)$
over $\{U_{1}(t),U_{2}(t)\}_{t=0}^{\infty},$
subject to $X(t+1)=\begin{bmatrix}F&G\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\begin{bmatrix}F^{\top}{}\\ G^{\top}\end{bmatrix},$
$\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\succeq{}0,\;X(0)=X_{0}\succeq{}0.$
In the above, the variables are the matrices $X(t)$, $U_{1}(t)$ and $U_{2}(t)$. The problem data is the matrices $S$, $R_{1}$, $R_{2}$, $F$ and $G$, where we additionally require that
$\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\succeq{}0.$ (7)
Although it may not be immediately clear at first sight, this is an optimization of precisely the type considered in 1($i$) where $\mathcal{P}$ is the positive semi-definite cone (which is positive, proper and self-dual).
We will now explain the correspondence between the variables and data in the above, and those in $1$The set of $p\times{}p$ positive semi-definite matrices is a proper cone in $\mathbb{R}^{p\left(p+1\right)/2}$, and so the mappings required to match the two problems explicitly can certainly be defined. However manipulations on the semi-definite cone are often more conveniently expressed in terms of matrices. We will have a need for such manipulations to identify the function $\phi$, and so proceed with this more informal connection between the two problems.. The variables are related as follows:
$x(t)$ $\longleftrightarrow{}X(t)$
$u(t)$ $\longleftrightarrow{}\begin{bmatrix}0&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}(t)\end{bmatrix}.$
That is the $x(t)$ and $u(t)$ variables correspond to a partitioning of a larger matrix. For the dynamics we have that:
$Ax(t)$ $\longleftrightarrow{}FX(t)F^{\top}$
$Bu(t)$ $\longleftrightarrow{}\begin{bmatrix}F&G\end{bmatrix}\begin{bmatrix}0&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\begin{bmatrix}F^{\top}{}\\ G^{\top}\end{bmatrix}.$
These dyanamics are clearly cone preserving under the constraint $\left(x,u\right)\in\mathcal{P}$, as required by 1. Similarly the costs can be connected through:
$s^{\top}x(t)$ $\longleftrightarrow{}\operatorname{tr}\left(SX(t)\right)$
$r^{\top}u(t)$ $\longleftrightarrow{}\operatorname{tr}\left(\begin{bmatrix}0&R_{1}\\ R_{1}^{\top}&R_{2}\end{bmatrix}\begin{bmatrix}0&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\right).$
Since the semi-definite cone is self-dual, under eq. 7 we have that $\left(s,r\right)\in\mathcal{P}^{*}$ as required. We have now established that our matrix optimization problem is on the form of 1($i$), and that all the prerequisits of the Theorem are satisfied except for the existence of the function $\phi$. We will now address this using 2.
To this end, let us now examine 2 when $\mathcal{P}$ is the semi-definite cone. Again it is helpful to relate vector and matrix variables, and so introduce
$\lambda$ $\longleftrightarrow{}\Lambda$
$\mu$ $\longleftrightarrow{}\begin{bmatrix}0&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}.$
After specialising the notation to this setting, 2 then states that the existence of $\phi\left(\mu\right)$ is equivalent to the set
$\mathcal{C}_{\mu}=\left\{\Lambda:\begin{bmatrix}\Lambda&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\succeq{}0\right\},$
having a minimum element (that is, an element $\Lambda^{*}\in\mathcal{C}_{\mu}$ such that $\Lambda^{*}\preceq{}\Lambda$ for all $\Lambda\in\mathcal{C}_{\mu}$). For the semi-definite cone whenever $\mathcal{C}_{\mu}$ is non-empty, this element exists, and is given by $\Lambda^{*}=M_{1}M_{2}^{+}M_{1}^{\top}{}$. Therefore
$\phi\left(\mu\right)\longleftrightarrow{}-M_{1}M_{2}^{+}M_{1}^{\top}{}.$
We are now ready to examine 1 ($ii$)–($iii$). Substituting in for all found correspondances, we see that in our matrix notation [^2] It is perhaps worth noting that from basic properties of adjoints, it follows that $A^{\top}\lambda(t)$ $\longleftrightarrow{}F^{\top}\Lambda(t)F$ $B^{\top}\lambda(t)$ $\longleftrightarrow{}\begin{bmatrix}0&F^{\top}\Lambda(t)G\\ G^{\top}\Lambda(t)F&G^{\top}\Lambda(t)G\end{bmatrix}.$ , the equation $\lambda_{*}=s+A^{\top}\lambda_{*}+\phi(r+B^{\top}\lambda_{*})$ becomes
$\Lambda^{*}=S+F^{\top}\Lambda^{*}F-\left(F^{\top}\Lambda^{*}G+R_{1}\right)\left(R_{2}+G^{\top}{}\Lambda^{*}G\right)^{+}\left(G^{\top}\Lambda^{*}F+R_{1}^{\top}{}\right).$
Note that a degree of care has to be taken when reading this formula, since the requirements on the domain of $\phi$ have been left implicit. For example, these are guaranteed to hold under the common assumption that $R_{2}\succ{}0$, after which the above collapses to the usual algebraic Riccati equation. This can then be solved in the usual way whenever the pair $\left(F,G\right)$ is stabilizable. 1 ($iii$) can be similarly analyzed. In this case, the convex program reduces to the semi-definite program
Maximize $\operatorname{tr}\left(\Lambda{}X_{0}\right)$
subject to $\begin{bmatrix}S-\Lambda&R_{1}\\ R_{1}^{\top}&R_{2}\end{bmatrix}+\begin{bmatrix}F^{\top}{}\\ G^{\top}{}\end{bmatrix}\Lambda\begin{bmatrix}F&G\end{bmatrix}\succeq{}0,\;\Lambda\succeq{}0.$
To conclude the example, let us finally connect explicitly to the standard LQR problem. The standard LQR problem is concerned with the optimization
Minimize $\sum_{t=0}^{\infty}\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}^{\top}\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}\text{ over }\{u(t)\}_{t=0}^{\infty}$
subject to $x(t+1)=Fx(t)+Gu(t),\,x(0)=x_{0}.$
To see the connection to the matrix optimization only requires us to notice that whenever
$\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}=\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}\begin{bmatrix}x(t)^{\top}&u(t)^{\top}\end{bmatrix},$ (8)
the specified matrix dynamics ensure that
$X(t+1)=x(t+1)x(t+1)^{\top}{},$
and also that
$\operatorname{tr}\left(\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\right)=$
$\qquad\qquad\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}^{\top}\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}.$
That is, under eq. 8, the costs and constraints in the standard LQR problem, and the matrix optimization problem, are identical. It then follows (with a little extra work) that the standard LQR problem corresponds to the special case of the matrix optimization when the initial condition $X_{0}$ is rank 1.
3.2 A Linear Regulator on a Polyhedral Cone
In this subsection we consider the application of 1 when $\mathcal{P}$ is the polyhedral cone (which are always positive and proper)
$\mathcal{P}=\left\{\begin{bmatrix}x\\ u\end{bmatrix}:\begin{bmatrix}I&0\\ E&-I\\ E&I\end{bmatrix}\begin{bmatrix}x\\ u\end{bmatrix}\geq{}0\right\},$
where $E\geq{}0$. The motivation for this choice comes from the study of positive systems. Note in particular that the condition $(x,u)\in\mathcal{P}$ constrains the system state to lie in the positive orthant, and the input to satisfy
$\left|{u}\right|\leq{}Ex.$
When minimising the cost in 1 ($i$), we are in effect optimizing performance under the constraint that the control keeps the system state positive. In what follows will recover the results of [5, Theorem 1].
First note that the dual cone $\mathcal{P}^{*}$ is also polyhedral, and is given by
$\mathcal{P}^{*}=\left\{\begin{bmatrix}\lambda\\ \mu\end{bmatrix}:\begin{bmatrix}\lambda\\ \mu\end{bmatrix}=\begin{bmatrix}I&0\\ E&-I\\ E&I\end{bmatrix}^{\top}\begin{bmatrix}w\\ y\\ z\end{bmatrix},\begin{bmatrix}w\\ y\\ z\end{bmatrix}\geq{}0\right\}.$ (9)
The conditions of 1 then provide conditions for solving the optimization in ($i$) whenever:
  1. 1)
    For all $\left(x,u\right)\in\mathcal{P}$, there exists a $v$ such that $\left(Ax+Bu,v\right)\in\mathcal{P}$.
  2. 2)
    $\left(s,r\right)\in\mathcal{P}^{*}$.
  3. 3)
    There exists a suitable function $\phi$.
It is readily checked that 1) amounts to requiring that $A\geq{}\left|{EB}\right|$. For 2) and 3) it is convenient to appeal to 2. Condition ($ii$) of this statement postulates that the set
$\mathcal{C}_{\mu}=\left\{\lambda:\left(\lambda,\mu\right)\in\mathcal{P}^{*}\right\}$
has a minimum element. From eq. 9 we see that for any $\mu$, $y$ and $z$ can be chosen according to
$y=\mu_{-}+v,\;z=\mu_{+}+v,$
where $v\geq{}0$, and the $\mu_{-}$ and $\mu_{+}$ denote the positive and negative parts of the vector $\mu$ respectively. It then follows that $\lambda\in\mathcal{C}_{\mu}$ if and only if
$\lambda=w+E^{\top}\left(\left|{\mu}\right|+v\right),$
where $w\geq{}0$. The minimum element is therefore given by $\lambda^{*}=E^{\top}\left|{\mu}\right|$. This implies that for (2) we require that $s\geq{}E^{\top}\left|{r}\right|$, and also that
$\phi\left(\mu\right)=-E^{\top}\left|{\mu}\right|.$
Condition ($ii$) in 1 then collapses to the existence of a $\lambda_{*}\geq{}0$ such that
$\lambda_{*}=s+A^{\top}\lambda_{*}-E^{\top}\left|{r+B^{\top}\lambda_{*}}\right|,$
which can be checked with linear programming. The convex program in ($iii$) similarly simplifies, this time to the following linear program:
Maximize $\lambda^{\top}x_{0}$
subject to $\begin{bmatrix}s\\ r\end{bmatrix}=\begin{bmatrix}I-A^{\top}&I&E^{\top}&E^{\top}\\ -B^{\top}&0&-I&-I\end{bmatrix}\begin{bmatrix}\lambda\\ w\\ y\\ z\end{bmatrix}\!,\begin{bmatrix}\lambda\\ w\\ y\\ z\end{bmatrix}\!\!\geq{}\!0.$
3.3 Linear Quadratic Regulation with Additional Structure
In this example we will continue our study of the matrix LQR problem outlined at the beginning of section 3.1. However, we now consider the case that the dynamics preserve a cone $\mathcal{P}$ that contains the semi-definite cone. When considering the application of 1, this will implicitly add further constraints to the allowable dynamics. However the upshot is that the optimal control will inherit these structural constraints, meaning that they can be exploited when implementing the optimal control laws. As we will see, this shows for example that the optimal control laws (under suitable definitions) for systems with a circulant structure inherit the same circulant structure [8, 9], with a host of other extensions, for example to systems defined through a generalized frequency variable [10].
The ideas that we are about to present can be significantly generalized, but should serve to illustrate the core ideas. We start with the definition of the cone.
Definition 1.
Let $Q\in\mathbb{R}^{m\times{}m}$ be invertible, and
$T=\begin{bmatrix}Q\otimes{}I_{p}&0\\ 0&Q\otimes{}I_{q}\end{bmatrix}.$
Define
$\mathcal{P}_{Q}=\left\{\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}{}&U_{2}\end{bmatrix}\in\mathbb{S}^{m\left(p+q\right)}:A\left(T\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}{}&U_{2}\end{bmatrix}T^{\top}{}\right)\succeq{}0\right\},$
where $\mathbb{S}^{m\left(p+q\right)}$ denotes the set of ${m\left(p+q\right)}\times{}{m\left(p+q\right)}$ symmetric matrices,
$A\left(\begin{bmatrix}W&X\\ Y&Z\end{bmatrix}\right)=\begin{bmatrix}\operatorname{diag}_{m}\left(W\right)&\operatorname{diag}_{m}\left(X\right)\\ \operatorname{diag}_{m}\left(Y\right)&\operatorname{diag}_{m}\left(Z\right)\end{bmatrix},$
and $\operatorname{diag}_{m}:\mathbb{R}^{mx\times{}my}\rightarrow{}\mathbb{R}^{mx\times{}my}$ denotes the operator that zeros out a given matrix outside a set of $m$ blocks of size $x\times y$ along the diagonal [^3] That is $\operatorname{diag}_{m}\left(M\right)=\begin{bmatrix}M_{11}&0&\cdots{}&0\\ 0&M_{22}&0\\ \vdots{}&0&\ddots{}&0\\ 0&\cdots{}&0&M_{mm}\end{bmatrix},M_{kk}\in\mathbb{R}^{x\times{}y}.$
We now have the following theorem, concerning the matrix LQR problem from section 3.1. The first part of the statement sets up the LQR problem for systems that preserve both the semi-definite cone, and $\mathcal{P}_{Q}$. The important part of the statement is eq. 10. In words, this equation shows that whenever the LQR problem is well defined, the optimal control can be computed through a state-feedback written in terms of matrices structured by the dual cone $\mathcal{P}_{Q}$, revealing the extra structure in the control law.
Theorem 3.
Let $\mathcal{P}_{Q}$ be as in 1, and suppose that $F$, $G$, $S$, $R_{1}$ and $R_{2}$ are such that:
  1. (a)
    $\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\in\mathcal{P}_{Q}$ implies that for some $V_{1}$ and $V_{2}$,
    $\begin{bmatrix}FXF^{\top}+FU_{1}G^{\top}+GU_{1}^{\top}F^{\top}+GU_{2}G^{\top}&V_{1}\\ V_{1}^{\top}&V_{2}\end{bmatrix}\in\mathcal{P}_{Q}.$
  2. (b)
    $\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}&R_{2}\end{bmatrix}\in\mathcal{P}_{Q}^{*}.$
Then whenever the optimization
Minimize $\sum_{t=0}^{\infty}\operatorname{tr}\left(\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\right)$
over $\{U_{1}(t),U_{2}(t)\}_{t=0}^{\infty},$
subject to $X(t+1)=\begin{bmatrix}F&G\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\begin{bmatrix}F^{\top}{}\\ G^{\top}\end{bmatrix},$
$\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\succeq{}0,X(0)\succeq{}0$
has a finite value, there exists a
$\begin{bmatrix}\Lambda{}&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}_{Q}^{*}$
such that the input sequence $\{\overline{U}_{1}(t),\overline{U}_{2}(t)\}_{t=0}^{\infty}$ that minimizes the cost satisfies
$\begin{bmatrix}\Lambda{}&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\begin{bmatrix}X(t)&\overline{U}_{1}(t)\\ \overline{U}_{1}(t)^{\top}&\overline{U}_{2}(t)\end{bmatrix}=0$ (10)
for all $t$.
Proof.
See Appendix A. ∎
To illustrate the theorem, suppose that the matrices $F$ and $G$ have a block-wise circulant structure. That is,
$F=\begin{bmatrix}f_{1}\bar{F}&f_{2}\bar{F}&\cdots{}&f_{m}\bar{F}\\ f_{m}\bar{F}&f_{1}\bar{F}&\cdots{}&f_{m-1}\bar{F}\\ \vdots{}&\vdots{}&\ddots{}&\vdots{}\\ f_{2}\bar{F}&f_{3}\bar{F}&\cdots{}&f_{1}\bar{F}\end{bmatrix},$
with a similar expression for $G$. We may write this compactly according to
$F=F_{\text{sd}}\otimes{}\bar{F}\;\;\text{and}\;\;G=G_{\text{sd}}\otimes{}\bar{G},$ (11)
where $\bar{F}\in\mathbb{R}^{p\times{}p}$ and $\bar{G}\in\mathbb{R}^{p\times{}q}$, and
$\!F_{\text{sd}}\!=\!\!\begin{bmatrix}f_{1}&f_{2}&\cdots{}&f_{m}\\ f_{m}&f_{1}&\cdots{}&f_{m-1}\\ \vdots{}&\vdots{}&\ddots{}&\vdots{}\\ f_{2}&f_{3}&\cdots{}&f_{1}\end{bmatrix}\!\!,G_{\text{sd}}\!=\!\!\begin{bmatrix}g_{1}&g_{2}&\cdots{}&g_{m}\\ g_{m}&g_{1}&\cdots{}&g_{m-1}\\ \vdots{}&\vdots{}&\ddots{}&\vdots{}\\ g_{2}&g_{3}&\cdots{}&g_{1}\end{bmatrix}\!\!.$
Critically for our purposes, the matrices $F_{\text{sd}}$ and $G_{\text{sd}}$ are simultaneously diagonalizable. That is, there exists an invertible matrix $Q$ such that both
$Q^{-1}F_{\text{sd}}Q\;\;\text{and}\;\;Q^{-1}G_{\text{sd}}Q$
are diagonal. This implies that
$\left(Q^{-1}\otimes{}I_{p}\right)F\left(Q\otimes{}I_{p}\right)=\operatorname{diag}_{m}\left(\left(Q^{-1}\otimes{}I_{p}\right)F\left(Q\otimes{}I_{p}\right)\right)$
and
$\left(Q^{-1}\otimes{}I_{p}\right)G\left(Q\otimes{}I_{q}\right)=\operatorname{diag}_{m}\left(\left(Q^{-1}\otimes{}I_{p}\right)G\left(Q\otimes{}I_{q}\right)\right).$
It follows from these expressions that for such matrices, 3(a) holds. As is shown in 1 in the Appendix A,
$\mathcal{P}_{Q}^{*}=\left\{T^{\top}A\left(Y\right)T:Y\succeq{}0\right\},$
where $A$ and $T$ are as in 1. That is, the dual cone consists of the positive semi-definite matrices with blocks that have a similar circulant structure to the matrices $F$ and $G$. 3 then states that provided the optimization criteria is chosen appropriately, the matrix that encodes the optimal control law eq. 10 will have the same block-wise circulant structure. Interestingly this cone based treatment makes no use of any orthogonality properties of the matrix $Q$ allowing similar insights to be derived for general dynamics that take the form of eq. 11, see for example [10].
4 Conclusions
We have proved a general result on optimal control on positive cones, covering standard linear quadratic control and corresponding results for positive systems as special cases. Of course, many other special cases can be derived by defining other relevant cones, or combinations of the ones above. Hopefully, this will in the future form the basis for a powerful toolbox combining the versatility of classical linear quadratic control with the rich scalability properties of positive systems and shortest path problems in networks. Preliminary steps in this direction have already been taken in [11] and [12].
Appendix A Proof of 3
The result will follow from an application of 1, based on the following simple properties of $\mathcal{P}_{Q}$.
Lemma 1.
Let $\mathcal{P}_{Q}$ be as in 1. Then:
  1. ($i$)
    If $P\succeq{}0$, then $P\in\mathcal{P}_{Q}$.
  2. ($ii$)
    $\begin{bmatrix}\Lambda&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}^{*}$
    only if
    $\begin{bmatrix}M_{1}M_{2}^{+}M_{1}^{\top}&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}^{*},$
    where $M_{2}\in\mathbb{R}^{mq\times{}mq}$.
Proof.
Claim ($i$) follows immediately from the implication
$Y\succeq{}0\implies{}A\left(Y\right)\succeq{}0.$
For ($ii$), we start by developing a formula for the dual cone. Let
$B\left(X\right)=A\left(TXT^{\top}\right).$
By Farkas’ lemma,
$\mathcal{P}^{*}=\left\{B^{*}\left(Y\right):Y\succeq{}0\right\}.$
Since for any symmetric matrices $C$ and $D$, $\operatorname{tr}\left(CD\right)=\operatorname{vec}\left(C\right)^{\top}\operatorname{vec}\left(D\right)$, we see that
$\operatorname{tr}\left(B\left(X\right)Y\right)$ $=\operatorname{tr}\left(A\left(TXT^{\top}{}\right)A\left(Y\right)\right)$ (12)
$=\operatorname{tr}\left(TXT^{\top}A\left(Y\right)\right)$
$=\operatorname{tr}\left(XT^{\top}A\left(Y\right)T\right).$
This shows that
$B^{*}\left(Y\right)=T^{\top}A\left(Y\right)T.$
It then follows that if
$\begin{bmatrix}\Lambda&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}^{*},$
then there exists a
$\begin{bmatrix}Y_{1}&Y_{2}\\ Y_{2}^{\top}&Y_{3}\end{bmatrix}=\begin{bmatrix}\operatorname{diag}_{m}\left(Y_{1}\right)&\operatorname{diag}_{m}\left(Y_{2}\right)\\ \operatorname{diag}\left(Y_{2}^{\top}\right)&\operatorname{diag}_{m}\left(Y_{3}\right)\end{bmatrix}\succeq{}0$
such that
$\begin{bmatrix}0&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}=T^{\top}\begin{bmatrix}0&Y_{2}\\ Y_{2}^{\top}&Y_{3}\end{bmatrix}T.$
It is then straightforward to show that
$\begin{bmatrix}M_{1}M_{2}^{+}M_{1}^{\top}&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}=T^{\top}\begin{bmatrix}Y_{2}Y_{3}^{+}Y_{2}^{\top}&Y_{2}\\ Y_{2}^{\top}&Y_{3}\end{bmatrix}T,$
which is in $\mathcal{P}^{*}$ as required. ∎
We now prove 3.
Proof.
Observe that by 1($i$), the optimization in 3 is lower bounded by the relaxed optimization problem
Minimize $\sum_{t=0}^{\infty}\operatorname{tr}\left(\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\right)$
over $\{U_{1}(t),U_{2}(t)\}_{t=0}^{\infty},$
subject to $X(t+1)=\begin{bmatrix}F&G\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\begin{bmatrix}F^{\top}{}\\ G^{\top}\end{bmatrix},$
$\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\in\mathcal{P}_{Q},X(0)\succeq{}0.$
An analogous argument to that made in section 3.1 and 1($ii$) then shows that the set
$\mathcal{C}_{\mu}=\left\{\Lambda:\begin{bmatrix}\Lambda&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}_{Q}^{*}\right\}$
is non-empty, it has a minimum element. Therefore under (a) and (b) in the theorem statement and by 2, the relaxed optimization problem meets all the conditions of 1, from which it soon follows that the relaxed optimization has a finite value if and only if there exists a
$\begin{bmatrix}\Lambda^{*}&0\\ 0&0\end{bmatrix}\in\mathcal{P}_{Q}^{*}$
such that
$\Lambda^{*}=S+F^{\top}\Lambda^{*}F-\left(F^{\top}\Lambda^{*}G+R_{1}\right)\left(R_{2}+G^{\top}{}\Lambda^{*}G\right)^{+}\left(G^{\top}\Lambda^{*}F+R_{1}^{\top}{}\right).$
Again by 1, the elements of the optimal input sequence $\overline{U}_{1}(t)$ and $\overline{U}_{2}(t)$ are the minimizers of the following optimization problem
Minimize $\operatorname{tr}\left(\begin{bmatrix}0&R_{1}+F^{\top}\Lambda^{*}G\\ R_{1}^{\top}+G^{\top}\Lambda^{*}F&R_{2}+G^{\top}\Lambda^{*}G\end{bmatrix}\begin{bmatrix}0&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\right)$
over $U_{1},U_{2},$
subject to $\begin{bmatrix}X(t)&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\in\mathcal{P}_{Q}.$
The result will then follow if
$\begin{bmatrix}X(t)&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\succeq{}0,$
since then all inputs to the relaxed problem are feasible also for the unrelaxed problem. It follows from an argument analogous to that in eq. 12, and from the assumption that $X(0)\succeq{}0$, that we may replace the constraint in the above with
$\begin{bmatrix}X(0)&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\succeq{}0$ (13)
without affecting the optimal value. Just as for the standard LQR problem it then follows from complementary slackness that there exists a $K$ such that
$\begin{bmatrix}I\\ K^{\top}\end{bmatrix}\Lambda^{*}\begin{bmatrix}I&K\end{bmatrix}\in\mathcal{P}_{Q}^{*}$
and
$\begin{bmatrix}I\\ K^{\top}\end{bmatrix}\Lambda^{*}\begin{bmatrix}I&K\end{bmatrix}\begin{bmatrix}X(0)&\overline{U}_{1}\\ \overline{U}_{1}^{\top}&\overline{U}_{2}\end{bmatrix}=0.$ (14)
Applying this control results in $X(1)\succeq{}0$, and rerunning this argument for all $t$ then shows that eqs. 13 and 14 hold for all $t$, not just $t=0$. Hence the optimal actions for the relaxed problem are feasible also for the unrelaxed problem, from which the result follows. ∎